OpenStudy (anonymous):

If 2x=3y=4z, then (x+y-z)/(x-y+z)=? @Hero

hero (hero):

This is a problem that is made to look difficult but it isn't.

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

@kryton1212 The 2x = 3y = 4z implies that each variable can be written in terms of any other variable. Hence you can re-write the expression in such a way that there is only 1 variable instead of 3.

hero (hero):

Wow, that was like a hit and run @genius12

OpenStudy (anonymous):

@genius12 how? 2/3=y/x 3/4=z/y ?

OpenStudy (anonymous):

@Hero Sorry but I didn't understand the metaphorical phrasing there.

hero (hero):

You posted your comment then ran away fast.

OpenStudy (anonymous):

But im still here..

hero (hero):

Weird, but okay

OpenStudy (agent0smith):

@kryton1212 not quite use 2x=3y=4z to get something like 2x = 3y which gives y = 2x/3 2x = 4z which gives z = 2x/4 Now you can replace y and z in the equation.

OpenStudy (anonymous):

are you kidding? 2x=2x=2x....

OpenStudy (agent0smith):

Huh? No. I don't know where you got that. If 2x=3y=4z then 2x=4z. \[\Large \frac{ x+y-z }{x-y+z } = \] \[\Huge \frac{ x+\frac{ 2x }{3 }-\frac{ x }{ 2 } }{x-\frac{ 2x }{3 }+\frac{ x }{ 2 } } = \]

OpenStudy (anonymous):

ok...

OpenStudy (agent0smith):

Now try simplifying that... just gotta add fractions, notice you can factor out the x\[\Huge \frac{x( 1+\frac{ 2 }{3 }-\frac{ 1 }{ 2 } )}{x(1-\frac{ 2}{3 }+\frac{ 1 }{ 2 } )} = \]

OpenStudy (anonymous):

\[\frac{ \frac{ 6x^{2}+x }{ 6x }}{ \frac{ 6x^{2}-6x }{ 6x } }\] seems complicated...

OpenStudy (agent0smith):

I don't know how you got there... see my last post.

OpenStudy (anonymous):

7/5 i think i am very stupid

OpenStudy (agent0smith):

Nope, 7/5 is correct :)

OpenStudy (anonymous):

thank you